# Straightening the Ruler: Field-Level Inference of the BAO Scale with LEFTfield

### v1

@@ Abstract
Current inferences of the BAO scale from galaxy clustering employ a reconstruction technique at fixed cosmology and bias parameters. Here, we present the first consistent joint Bayesian inference of the isotropic BAO scale, jointly varying the initial conditions as well as all bias coefficients, based on the EFT-based field-level forward model LEFTfield. We apply this analysis to mock data generated with a much higher cutoff (resolution), resulting in a significant model mismatch between mock data and the model used in the inference. We demonstrate that the remaining systematic bias in the BAO scale is below 2% for all data considered and below 1% when Eulerian bias is used for inference. Furthermore, we find that the inferred error on the BAO scale is up to 1.6 times smaller compared to that from a replication of the standard power-spectrum reconstruction approach, using the same scales as in the field-level inference, with the improvement growing towards smaller scales (higher $k$). Thus, a field-level approach to BAO not only allows for a consistent inference of the BAO scale, but promises to achieve more precise measurements on the same data as well.
@@ 1 Introduction
Baryon Acoustic Oscillations (BAO) represent one of the most significant observables in cosmology. Originating from the period when photons and baryons were tightly coupled, the imprints of BAO are visible in both the Cosmic Microwave Background (CMB) temperature anisotropies and the clustering of matter at later times. By measuring the apparent size of the BAO scale, $r_{s}$, in the late-time matter distribution, it is possible to estimate both the angular diameter distance and the Hubble parameter as functions of redshift. Moreover, the BAO feature imprinted in galaxy clustering is much less sensitive to the details of the galaxy-matter relation (bias) than the broad-band galaxy statistics. However, the nonlinear evolution of the matter density field and the formation of structures complicate this picture. The bulk motion of matter (and galaxies with it) leads to a damping of the BAO oscillations in the power spectrum, reducing the accuracy with which the BAO scale can be measured from the late-time clustering of galaxies [1, 2, 3, 4].
@@
Traditional BAO inference methods attempt to correct for the bulk flows by performing a nonlinear transformation of the data. Central to this is the estimation of displacements from the observed density field to move galaxies to their estimated initial positions [5, 6, 7, 8, 9, 10, 11, 12, 13, 14]. This method does require making assumptions about cosmology and bias, and for a consistent inference one should really sample the parameters associated with these together with the BAO scale. Here, we aim to deal with this issue by adopting a fully Bayesian field-level forward modeling approach. This approach incorporates BAO reconstruction by jointly inferring the large-scale modes responsible for the bulk flow [15]. Crucially, it does not rely on a fiducial cosmology, as usually assumed in the reconstruction procedure. Instead, cosmological parameters are jointly and consistently inferred with the initial conditions and bias parameters. [^1] Note that in this work, we also fix all cosmological parameters apart from the BAO scale in the inference, but all bias parameters including the linear bias $b_{1}$ are jointly inferred with the BAO scale. Furthermore, since the field-level approach does not rely on compression functions, but instead obtains the information directly from the field, it also extracts the BAO information contained in higher $n$-point functions such as the bispectrum.
@@
We have already discussed these issues in our previous work [16] where the field-level approach for inferring the BAO scale was first introduced. The starting point is the joint posterior for the initial density field, cosmological parameters, and nuisance parameters, which include bias parameters and stochastic amplitudes. At the heart of the field-level posterior lies the likelihood function, which is the key ingredient allowing us to access the information at the field level. The likelihood we use was derived within the framework of the effective field theory (EFT) of large scale structure (LSS) [17, 18] in a series of papers [15, 19, 20] and we will refer to it as the EFT likelihood or field-level likelihood. In every EFT theory, a fundamental component is the cutoff scale, denoted as $\Lambda$, representing the maximum wavenumber of modes considered in the calculations (in $n$-point-function-based analyses, this is often denoted with $k_{\rm max}$). While the specific value of $\Lambda$ is arbitrary, the upper limit for $\Lambda$ in the case of the EFT of LSS is the nonlinearity scale ($\Lambda_{\text{NL}}\approx 0.25\,h\,{\rm Mpc}^{-1}$ at $z=0$), where perturbation theory of LSS breaks down. In addition to the BAO scale, EFT-based field-level inference results have been applied to the $\sigma_{8}$ parameter in [21, 22, 23, 24] as well as the growth rate $f$ in [25].
@@
In our previous paper on BAO [16], we analyzed dark matter halo samples in N-body simulations, but in the scenario of fixed initial conditions. To assess the effectiveness of the field-level approach combined with the EFT likelihood, we compared it to a power spectrum likelihood whose covariance was derived to account for the fixed initial conditions (i.e., no cosmic variance). Our analysis revealed that the error bars for the BAO scale were between 1.1 and 3.3 times smaller when using the field-level likelihood as compared to the power spectrum without BAO reconstruction, where this range corresponds to the range of scales considered, $\Lambda=[0.1-0.25]\,h\,\text{Mpc}^{-1}$, as well as different halo samples and redshifts.
@@
In this paper, we go a significant step further and use the field-level likelihood to perform a joint inference of the BAO scale and the initial conditions, which of course is necessary in the application to real data. Sampling initial conditions is challenging since the dimensionality of the parameter space is of the order of $10^{5}-10^{6}$. To deal with this challenge, we follow the lead of [26] and use the Hamiltonian Monte Carlo (HMC) [27] sampling method to sample the initial conditions and slice sampler [28] to sample the cosmological parameters. We apply this analysis to mock data generated with the LEFTfield code, a Lagrangian EFT-based forward model [29], which we also use for the inference. Importantly, our mock data are generated with a substantially higher cutoff than that used in the inference, as well as different bias models (with bias parameters representative of actual halo samples), so that we can test the robustness against model mis-specifications.
@@
In a pioneering work, Ref. [30] demonstrated field-level inference of cosmology and distances on mock data, obtaining very tight constraints on the expansion history. Differently to this work, however, they did not separate BAO and broad-band (Alcock-Paczyński distortion) contributions. In addition, our forward model differs substantially, in that we use an EFT-based forward model, and, crucially, our results also incorporate model mis-specification.
@@
In Section 2, we briefly introduce the forward model used in this paper. Section 3 provides details of the implementation within the LEFTfield code. Section 4 elaborates on the generation of mock data. Finally, in Section 5.2, we present the results from the joint inference of BAO along with the initial conditions, and compare to the power-spectrum analysis with and without reconstruction applied in Section 6.
@@ 2 Forward Model
The forward model employed here was introduced in [15]. The main object of this model is a joint posterior for the initial density field $\delta_{\rm in}$, cosmological parameters $\theta$ and “nuisance” parameters (bias parameters {$b_{O}$} and stochastic amplitudes $\sigma$) given the data, i.e. $P(\delta_{\rm in},\theta,\{b_{O}\},\sigma|\delta_{g})$. There are four ingredients to this posterior:
@@
$\delta_{\rm in}(\boldsymbol{k},\hat{s})=W_{\Lambda}(k)\sqrt{P_{\text{L}}(k)}\hat{s}(\boldsymbol{k}),$ (2.2)
where $W_{\Lambda}(k)$ is a sharp-$k$ filter which we use to ensure the proper renormalization of the evolution of large-scale modes, $P_{\rm L}$ is the linear power spectrum and $\hat{s}(\boldsymbol{k})$ is the unit Gaussian field which we use to represent the normalized initial conditions (sometimes loosely referred to as “phases”).
@@
To achieve our goal, we start by introducing the parameter $\beta$ in the standard linear power spectrum approximation,
$P_{\rm L}(k,\beta)=P_{\rm L,sm}(k)[1+A\sin(k\,\beta\,r_{\rm fid}+\phi)\exp(-k/k_{\rm D})].$ (2.3)
Here, $P_{\rm L,sm}(k)$, $A$, $\phi$, and $k_{\rm D}$ are determined by fitting of Eq. (2.3) to the linear power spectrum obtained using the CLASS code [33]. The parameter $r_{\rm fid}$ signifies the fiducial BAO scale. Additionally, $\beta$ is defined as the ratio of a proposed BAO scale size $r_{s}$ to its fiducial size,
@@
$\delta_{\beta}(k,\beta)=f(k,\beta)\delta_{\rm fid}(k).$ (2.6)
Therefore, the fields $\delta_{\beta}(k,\beta)$ and $\delta_{\rm fid}(k)$ differ only in the size of the BAO scale; all other parameters are set to the exact same values.
@@
In this paper, we use the Lagrangian perturbation theory (LPT). If we assume that at the initial time $\tau=0$ the particle was at the position q, which we call Lagrangian position, then under the gravitational evolution it will move and we will finally observe it at the Eulerian position x at some time $\tau$. The vector field connecting those two positions is called the Lagrangian displacement field $\boldsymbol{\psi}(\textbf{q},\tau)$ and it is the main object of LPT,
$\textbf{x}(\textbf{q},\tau)=\textbf{q}+\boldsymbol{\psi}(\textbf{q},\tau).$ (2.7)
@@
$\boldsymbol{\psi}(\textbf{q},\tau)=\sum_{n=1}^{\infty}\boldsymbol{\psi}^{(n)}(\textbf{q},\tau),$ (2.8)
@@
One of the key advantages of LPT lies in the possibility to integrate the equations of motion for the displacement field analytically, yielding recursion relations for $\boldsymbol{\psi}^{(n)}$ as a function of $\boldsymbol{\psi}^{(m)}$ with $m<n$. This feature enables a relatively straightforward construction of the displacement field [34, 35, 36, 37, 29] (see [38] for a related construction based on Eulerian standard perturbation theory). The starting point of these recursion relations is given by
$\cdot\boldsymbol{\psi}^{(1)}(\textbf{q},\tau)=-\delta^{(1)}(\textbf{q},\tau)=-D_{\rm norm}(\tau)\delta_{\rm in}(\textbf{q}),$ q ) , (2.9)
where $D_{\rm norm}$ is the normalized growth factor relative to the epoch of the linear power spectrum used in the definition of the initial conditions, Eq. (2.2). The matter density field is given fully nonlinearly by $1+\delta(x,\tau)=|\textbf{1}+\boldsymbol{\nabla}_{q}\boldsymbol{\psi}(q)|^{-1}$. We do not need to evaluate this equation explicitly however, since we instead obtain $\delta$ by assigning pseudo-particles to a grid at their Eulerian positions given by Eq. (2.7) (see Sec. 3).
@@
To find a biased tracer density field, we use the deterministic relation
@@
In the case of the Eulerian bias expansion, bias operators are constructed from the matter density field obtained using the LPT as described in the previous section,
@@
We also include the higher derivative term $\nabla^{2}\delta$ which accounts for the non-locality of halo formation. The coefficients of higher-derivative operators are related to the spatial scale $R_{*}$ which quantifies the size of the spatial region involved in the process of halo formation, and their contribution is suppressed by $k^{2}R_{*}^{2}$ on large scales.
@@
the operators $O^{L}$ are constructed from all scalar contributions to the symmetric part of the tensor $\textbf{M}(\textbf{q},\tau)$, defined as
$M_{ij}(\textbf{q},\tau)\equiv\partial_{q,i}\psi_{j}(\textbf{q},\tau).$ (2.15)
We do not need to include $\text{tr}[M^{(n)}]$ for $n>1$, as they can always be expressed in terms of scalars constructed using lower-order operators [37]. Here, we again focus solely on the second-order bias expansion. Therefore, the bias operators we use are [40]
@@
By construction, our model does not capture modes above the cutoff $\Lambda$ in the initial conditions. These modes introduce scatter around the mean relation $\delta_{g,\rm det}$, which is described by the stochastic contribution expressed as the field $\epsilon$ in Eq. (2.1). The purpose of the field-level EFT likelihood is to absorb this stochastic contribution. Considering that $\epsilon$ originates from integrating out numerous independent $k$ modes, the central limit theorem ensures that this field is Gaussian at leading order. Furthermore, owing to the local nature of tracer formation, the power spectrum of $\epsilon$ is constant at leading order, with corrections scaling as $k^{2}$. These properties lead to the formulation of the EFT likelihood
$\ln\mathcal{L}_{\text{EFT}}(\delta_{g}|\hat{s},\theta,\{b_{O}\},\sigma)=-\frac{1}{2}\sum_{k<\Lambda}\Bigg{[}\ln[2\pi\sigma^{2}(k)]+\frac{1}{\sigma^{2}(k)}|\delta_{g}(\textbf{k})-\delta_{g,{\rm det}}[\hat{s},\theta,\{b_{O}\}](\textbf{k})|^{2}\Bigg{]}.$ (2.17)
@@
$\sigma^{2}(k)=\sigma_{\epsilon}^{2}(k)(1+k^{2}\sigma_{\epsilon,2})^{2}.$ (2.18)
The parameter $\sigma_{\epsilon}$ is the variance of the noise field $\epsilon$ on the discrete grid. This quantity is dependent on both the grid size, $N_{g}^{\Lambda}$, and the box size $L$. Its relationship with the noise power spectrum $P_{\epsilon}$ is then given by

### v2

@@ Abstract
Current inferences of the BAO scale from galaxy clustering employ a reconstruction technique at fixed cosmology and bias parameters. Here, we present the first consistent joint Bayesian inference of the isotropic BAO scale, jointly varying the initial conditions as well as all bias coefficients, based on the EFT-based field-level forward model LEFTfield. We apply this analysis to mock data generated at a much higher cutoff, or resolution, resulting in a significant model mismatch between mock data and the model used in the inference. We demonstrate that the remaining systematic bias in the BAO scale is below 2% for all data considered and below 1% when Eulerian bias is used for inference. Furthermore, we find that the inferred error on the BAO scale is typically 30%, and up to 50%, smaller compared to that from a replication of the standard post-reconstruction power-spectrum approach, using the same scales as in the field-level inference. The improvement in BAO scale precision grows towards smaller scales (higher $k$). As a validation test, we repeat this comparison on a mock dataset that is linearly biased with respect to a 1LPT (Zel’dovich) density field, following the assumption made in standard reconstruction approaches. We find that field-level inference indeed yields the same error bar as the post-reconstruction power spectrum, which is expectd to be optimal in this case. In summary, a field-level approach to BAO not only allows for a consistent inference of the BAO scale, but promises to achieve more precise measurements on realistic, nonlinearly biased tracers as well.
@@ 1 Introduction
Baryon Acoustic Oscillations (BAO) represent one of the most significant observables in cosmology. Originating from the period when photons and baryons were tightly coupled, the imprints of BAO are visible in both the Cosmic Microwave Background (CMB) temperature anisotropies and the clustering of matter at later times. By measuring the apparent size of the BAO scale, $r_{s}$, in the late-time matter distribution, it is possible to estimate both the angular diameter distance and the Hubble parameter as functions of redshift. Moreover, the BAO feature imprinted in galaxy clustering is much less sensitive to the details of the galaxy-matter relation (bias) than the broad-band galaxy statistics. However, the nonlinear evolution of the matter density field and the formation of structures does have an impact. The bulk motion of matter (and galaxies with it) leads to a damping of the BAO oscillations in the power spectrum, reducing the accuracy with which the BAO scale can be measured from the late-time clustering of galaxies, as well as inducing a small systematic shift [1, 2, 3, 4].
@@
Traditional BAO inference methods attempt to correct for the bulk flows by performing a nonlinear transformation of the data, known as BAO reconstruction. Central to this is the estimation of displacements from the observed galaxy density field to move galaxies to their estimated initial positions [5, 6, 7, 8, 9, 10, 11, 12, 13, 14]. This method however requires making assumptions about cosmology and bias; in a consistent inference framework one should really sample the parameters associated with these together with the BAO scale. Here, we aim to deal with this issue by adopting a fully Bayesian field-level forward modeling approach. This approach incorporates BAO reconstruction by jointly inferring the large-scale modes responsible for the bulk flow [15]. Crucially, it does not rely on a fiducial cosmology, as usually assumed in the reconstruction procedure. Instead, cosmological parameters are jointly and consistently inferred with the initial conditions and bias parameters. [^1] Note that in this work, we also fix all cosmological parameters apart from the BAO scale in the inference, but all bias parameters including the linear bias $b_{1}$ are jointly inferred with the BAO scale. Furthermore, since the field-level approach does not rely on compression functions, but instead obtains the information directly from the field, it also extracts the BAO information contained in higher $n$-point functions such as the bispectrum, as opposed to the standard procedure that only uses the reconstructed power-spectrum information.
@@
We have already discussed these issues in our previous work [16] where the field-level approach for inferring the BAO scale was first introduced. The starting point is an expression for the joint posterior for the initial density field, cosmological parameters, and other model parameters, which include bias parameters and stochastic amplitudes, derived within the framework of the effective field theory (EFT) of large-scale structure (LSS) [17, 18]. At the heart of the field-level posterior lies the likelihood function, which is the key ingredient allowing us to access the information at the field level. The likelihood was derived within in a series of papers [15, 19, 20] and we will refer to it as the EFT likelihood or field-level likelihood. In every EFT theory, a fundamental component is the cutoff scale, denoted as $\Lambda$, representing the maximum wavenumber of modes considered in the calculations (in $n$-point-function-based analyses, this is often denoted with $k_{\rm max}$). While the specific value of $\Lambda$ is arbitrary, the upper limit for $\Lambda$ in the case of the EFT of LSS is the nonlinearity scale ($\Lambda_{\text{NL}}\approx 0.25\,h\,{\rm Mpc}^{-1}$ at $z=0$), where perturbation theory of LSS breaks down.
The results presented here are obtained with the LEFTfield code, a Lagrangian EFT-based forward model [21]. Previous cosmology results with LEFTfield have been presented in [22, 23, 24, 25] for the amplitude of perturbations ($\mathcal{A}_{s}$ or equivalently $\sigma_{8}$), as well as the growth rate $f$ in [26]. Moreover, Ref. [27] (“Beyond-2pt challenge”) recently presented the results of a variety of inference methods applied to cosmology-blinded mock catalogs constructed by populating dark matter halos with a halo occupation distribution, where field-level inference based on LEFTfield showed very competitive results for $\sigma_{8}$ inference.
@@
In our previous paper on BAO [16], we inferred the BAO scale on dark matter halo samples in N-body simulations with LEFTfield, but in the scenario of fixed initial conditions. To assess the effectiveness of the field-level approach combined with the EFT likelihood, we compared it to a power spectrum likelihood whose covariance takes into account the fixed initial conditions (i.e., no cosmic variance). Our analysis revealed that the error bars for the BAO scale were between 1.1 and 3.3 times smaller when using the field-level likelihood as compared to the power spectrum without BAO reconstruction, where this range corresponds to the range of scales considered, $\Lambda=[0.1-0.25]\,h\,\text{Mpc}^{-1}$, as well as different halo samples and redshifts.
@@
In this paper, we go a significant step further and use the field-level likelihood to perform a joint inference of the BAO scale and the initial conditions, which of course is necessary in the application to real data. Sampling initial conditions is challenging since the dimensionality of the parameter space is of the order of $10^{6}$. To deal with this challenge, we follow the lead of [28] and use the Hamiltonian Monte Carlo (HMC) [29] sampling method to sample the initial conditions, and slice sampler [30] to sample the cosmological parameters. We apply this analysis to mock data generated with LEFTfield. Importantly, our mock data are generated with a substantially higher cutoff than that used in the inference, as well as different bias models (with bias parameters representative of actual halo samples), so that we can test the robustness against model misspecification.
@@
In a pioneering work, Ref. [31] demonstrated field-level inference of cosmology and distances on mock data, obtaining very tight constraints on the expansion history. Differently to this work, however, they did not separate BAO and broad-band (Alcock-Paczyński distortion) contributions. In addition, our forward model differs substantially, in that we use an EFT-based forward model, and, crucially, our results also incorporate model misspecification.
@@
In Section 2, we briefly introduce the forward model used in this paper. Section 3 provides details of the implementation within the LEFTfield code. Section 4 elaborates on the generation of mock data. In Section 5, we then present results from the joint inference of BAO along with the initial conditions, beginning with several validation tests of the inference pipelines. We describe the BAO reconstruction procedure, and compare field-level inference to the pre- and post-reconstruction power-spectrum analysis in Section 6. The appendix contains MCMC diagnostics and further details on the field-level inference results.
@@ 2 Forward Model
The forward model employed here was introduced in [15]. The main object of this model is a joint posterior for the initial density field $\delta_{\rm in}$, cosmological parameters $\theta$ and “nuisance” parameters (bias parameters {$b_{O}$} and stochastic amplitudes $\sigma$) given the data, i.e., $P(\delta_{\rm in},\theta,\{b_{O}\},\sigma|\delta_{g})$. There are four ingredients to this posterior:
@@
$\delta_{\rm in}(\bm{k},\hat{s})=W_{\Lambda}(k)\sqrt{P_{\text{L}}(k)}\hat{s}(\bm{k}),$ (2.2)
where $W_{\Lambda}(k)$ is a sharp-$k$ filter which we use to ensure the proper renormalization of the evolution of large-scale modes, $P_{\rm L}$ is the linear power spectrum and $\hat{s}(\bm{k})$ is the unit Gaussian field which we use to represent the normalized initial conditions (sometimes loosely referred to as “phases”).
@@
To achieve this, we parametrize the linear power spectrum via
$P_{\rm L}(k,\beta)=P_{\rm L,sm}(k)[1+A\sin(k\,\beta\,r_{\rm fid}+\phi)\exp(-k/k_{\rm D})],$ (2.3)
where $P_{\rm L,sm}(k)$, $A$, $\phi$, and $k_{\rm D}$ are determined by fitting Eq. (2.3) to the linear power spectrum obtained using the CLASS code [34] for the fiducial cosmology. The parameter $r_{\rm fid}$ signifies the fiducial BAO scale, and $\beta$ is defined as the ratio of a proposed BAO scale size $r_{s}$ to its fiducial size,
@@
$\delta_{\beta}(\textbf{k},\beta)=f(k,\beta)\delta_{\rm fid}(\textbf{k}).$ (2.6)
Therefore, the fields $\delta_{\beta}(\textbf{k},\beta)$ and $\delta_{\rm fid}(\textbf{k})$ differ only in the size of the BAO scale; all other parameters are kept fixed to the same values.
@@
In this paper, we use the Lagrangian perturbation theory (LPT). If we assume that at the initial time $\tau=0$ the particle was at the comoving position q, which we call Lagrangian position, then under the gravitational evolution it will move and we will finally observe it at the comoving Eulerian position x at some time $\tau$. The vector field connecting those two positions is called the Lagrangian displacement field $\bm{\psi}(\textbf{q},\tau)$ and it is the main object of LPT,
$\textbf{x}(\textbf{q},\tau)=\textbf{q}+\bm{\psi}(\textbf{q},\tau).$ (2.7)
@@
$\bm{\psi}(\textbf{q},\tau)=\sum_{n=1}^{\infty}\bm{\psi}^{(n)}(\textbf{q},\tau),$ (2.8)
@@
One of the key advantages of LPT lies in the possibility to integrate the equations of motion for the displacement field analytically, yielding recursion relations for $\bm{\psi}^{(n)}$ as a function of $\bm{\psi}^{(m)}$ with $m<n$. This feature enables a relatively straightforward construction of the displacement field [35, 36, 37, 38, 21] (see [39] for a related construction based on Eulerian standard perturbation theory). The starting point of these recursion relations is given by
$\cdot\bm{\psi}^{(1)}(\textbf{q},\tau)=-\delta^{(1)}(\textbf{q},\tau)=-D_{\rm norm}(\tau)\delta_{\rm in}(\textbf{q}),$ q ) , (2.9)
where $D_{\rm norm}$ is the normalized growth factor relative to the epoch of the linear power spectrum used in the definition of the initial conditions, Eq. (2.2). The matter density field is given fully nonlinearly by $1+\delta(x,\tau)=|\textbf{1}+\bm{\nabla}_{q}\bm{\psi}(q)|^{-1}$. We do not need to evaluate this equation explicitly however, since we instead obtain $\delta$ by assigning pseudo-particles to a grid at their Eulerian positions given by Eq. (2.7) (see Sec. 3).
@@
To predict the biased tracer density field, we use the deterministic relation
@@
In the case of the Eulerian bias expansion, bias operators are constructed from the matter density field obtained using LPT as described in the previous section,
@@
We also include the higher-derivative term $\nabla^{2}\delta$ which accounts for the non-locality of halo formation. The coefficients of higher-derivative operators are related to the spatial scale $R_{*}$ which quantifies the size of the spatial region involved in the process of halo or galaxy formation, and its contribution is suppressed by $k^{2}R_{*}^{2}$ on large scales.
@@
the operators $O^{L}$ are constructed from all scalar contributions to the symmetric part of the Lagrangian distortion tensor, defined as
$M_{ij}(\textbf{q},\tau)\equiv\partial_{q,(i}\psi_{j)}(\textbf{q},\tau).$ (2.15)
We do not need to include $\text{tr}[M^{(n)}]$ for $n>1$, as these terms can always be expressed in terms of scalars constructed using lower-order operators [38]. Here, we again focus solely on the second-order bias expansion. Therefore, the bias operators we use are [41]
@@
By construction, our model does not capture modes above the cutoff $\Lambda$ in the initial conditions. These modes introduce scatter around the mean relation $\delta_{g,\rm det}$, which is described by the stochastic contribution expressed as the field $\epsilon$ in Eq. (2.1). The purpose of the field-level EFT likelihood is to absorb this stochastic contribution. Considering that $\epsilon$ originates from integrating out numerous independent $k$ modes, the central limit theorem ensures that this field is Gaussian at leading order. Furthermore, owing to the local nature of tracer formation, the power spectrum of $\epsilon$ is constant at leading order, with higher-derivative corrections scaling as $k^{2}$. These properties lead to the formulation of the EFT likelihood
$\ln\mathcal{L}_{\text{EFT}}(\delta_{g}|\hat{s},\theta,\{b_{O}\},\sigma)=-\frac{1}{2}\sum_{|\textbf{k}|\leqslant\Lambda}\Bigg{[}\ln[2\pi\sigma^{2}(k)]+\frac{1}{\sigma^{2}(k)}|\delta_{g}(\textbf{k})-\delta_{g,{\rm det}}[\hat{s},\theta,\{b_{O}\}](\textbf{k})|^{2}\Bigg{]}.$ (2.17)
@@
$\sigma^{2}(k)=\sigma_{\epsilon}^{2}\>(1+k^{2}\sigma_{\epsilon,2})^{2}.$ (2.18)
The parameter $\sigma_{\epsilon}^{2}$ quantifies the variance of the noise field $\epsilon$ on the discrete grid. Its relationship with the noise power spectrum in the low-$k$ limit, $P_{\epsilon}$, is given by [24]